A piece of resistant wire of diametre of 0.8mm and resistant is 6ohm's has a resistivity of 8.8*10^4 metre.calculate the length of the wire take π=22/7


1
Expert's answer
2021-11-19T10:52:03-0500

R"= \rho\frac{l}{A}"


Where "R=" Resistance

"\rho" "=" Resistivity

L "=" Length

A "=" Cross section area "=\pi\>r^2"




"A=\frac{22}{7}(\frac{0.8}{2}×10^{-3})^2"


"=5.029\u00d710^{-7}m^2"



"l=\frac{AR}{\rho}"


NB: Value of resistivity given in the question is

"\color{blue}{8.8×10^4metres}"

1. The units are wrong

2 The value is unrealistically high for a wire


Taking the value of resistivity as :

"\textcolor{red}{8.8×10^{-4}\>\Omega m}"


"Length=\frac{5.029×10^{-7}×6}{8.8×10^{-4}}"



"=3.429\u00d710^{-3}m"